# Rigidity aspects of a cosmological singularity theorem

### v2

@@ Abstract
Improving a singularity theorem in General Relativity by Galloway and Ling we show the following (cf. Theorem 1): If a globally hyperbolic spacetime $M$ satisfying the null energy condition contains a closed, spacelike Cauchy surface $(V,g,\cal K)$ (with metric $g$ and extrinsic curvature $\cal K$}) which is 2-convex (meaning that the sum of the lowest two eigenvalues of $\cal K$} is non-negative), then either $M$ is past null geodesically incomplete, or $V$ is a spherical space, or $V$ or some finite cover $\tilde{V}$ is a surface bundle over the circle, with totally geodesic fibers. Moreover, (cf. Theorem 2) if $(V,g,{\cal K})$ admits a $U(1)$ isometry group with corresponding Killing vector $\xi$, we can relax the convexity requirement in terms of a decomposition of $\cal K$} with respect to the directions parallel and orthogonal to $\xi$. Finally, (cf. Propositions 1-3) in the special cases that $V$ is either non-orientable, or non-prime, or an orientable Haken manifold with vanishing second homology, we obtain stronger statements in both Theorems without passing to covers.
@@ 1 Introduction
The classic singularity theorems due to Hawking and Penrose [1] yield geodesic incompleteness in different settings, namely "cosmological" and "black hole" ones. In the former setting one requires a compact Cauchy surface and a condition on its second fundamental form, while the latter setting assumes non-compact data containing a trapped or marginally trapped surface. In either case, convergence ("energy-") conditions on the spacetime Ricci tensor $\mathrm{Ric}$ are also necessary. In the sequel we will only need the null energy condition (NEC), namely $\mathrm{Ric}(X,X)\geq 0$ for all null vectors $X$.
@@
    Assume $V$ is $2$-convex, i.e. the sum of the smallest two eigenvalues of the future second fundamental form $\mathcal{K}$ is nonnegative.
@@
We conclude here with sketching the strategy of our proofs. For simplification, we assume that $V$ is not a spherical space, and we pass to a finite, orientable cover $\tilde{V}$ whenever necessary or useful. Our key prerequisite is the prime decomposition of orientable 3-manifolds, see (3.1) below. Moreover, from the positive resolution of the virtual positive $b_{1}$-conjecture [5], it follows that $V$ or a finite cover $\tilde{V}$ has vanishing second homology $H_{2}$. As is well-known, this implies that $V$ (or $\tilde{V}$) contains an embedded, non-separating, two-sided minimal surface $\Sigma$ of least area. By a known Lemma (Lemma 0), it follows that a neighborhood of $\Sigma$ can be foliated by CMC surfaces $\Sigma_{t}$. To prove Theorem 1 we now invoke the convexity requirement on $\cal K$} which, via (1.1), yields control on the signs of the null expansions $\theta^{\pm}$ on each $\Sigma_{t}$. Together with the null energy condition, this allows an application of Penrose’s singularity theorem on a suitable infinite cover of $V$ (or $\tilde{V}$). This gives past null geodesic incompleteness except for the degenerate cases stated in (ii), (iii) of Theorem 1. In case (iii) compactness theorems allow us to extend the local folitaion of Lemma 1 to all of $V$.
@@
We remark that Theorem 1 can be proven in a different way, namely by using the positive resolution of the virtual Haken conjecture and Proposition 1. In fact, this detour was taken by the authors originally. We are grateful to Greg Galloway for bringing to our attention the positive resolution of the virtual positive $b_{1}$ conjecture which greatly simplifies the proof. A similar simplification also applies to the proof of the main result in [2], reformulated as Theorem 0 above.
@@ 2 Preliminaries
The following lemma is a standard application of the inverse function theorem. For a proof of the more involved $\lambda=0$ case, see the proof of Thm. 2.38 in [6] or [3, Appendix A].
@@
Let $\Sigma$ be an embedded stable minimal surface within a Riemannian manifold $V$. Then there is a neighborhood $U$ of $\Sigma$ within $V$, diffeomorphic to $(-\epsilon,\epsilon)\times\leavevmode\nobreak\ \Sigma$, such that each leaf $\Sigma_{t}:=\{t\}\times\Sigma$ has constant mean curvature.
@@
Proof. The following facts will be used to prove this Lemma as well as Lemma 2. First, there is a spacetime cover $P\colon\tilde{M}\to M$ such that $p=P|_{\tilde{V}}$, $\tilde{V}$ is a Cauchy surface for $\tilde{M}$, and the second fundamental form on $\tilde{V}$ is the pullback $\tilde{\cal K}=P^{*}\cal K$, see [2, Lem. 4]. Moreover, the null convergence condition also “lifts”, and $M$ is null geodesically complete if and only if $\tilde{M}$ is. Lastly, if a surface $\Sigma^{\prime}$ embedded in $V$ is constructed via a foliation from $\Sigma$ within $V$, then $\Sigma^{\prime}$ lifts to some $\tilde{\Sigma}^{\prime}$, which, like $\tilde{\Sigma}$, also separates $\tilde{V}$ into two noncompact ends.
@@
2.) $\mathbf{H_{t}\equiv 0\leavevmode\nobreak\ \forall\leavevmode\nobreak\ t}$.
@@
say $\tilde{W}_{t}^{+}$, does not vanish identically on $\tilde{\Sigma}_{t}$. By virtue of Raychaudhuri’s equation and another application of Lemma 5.2 in [8] we find surfaces $\tilde{\Sigma}^{\prime}$ near the past light cone of $\tilde{\Sigma}_{t}$ with $\tilde{\theta}^{\prime+}<0$. Another application of Penrose’s theorem in the cover $\tilde{M}$ and projecting down to $M$ contradicts past null geodesic completeness within $M$. Therefore both $(+)$-terms on the r.h.s. of (2.2) have to vanish individually. Clearly, the case $\tilde{W}^{-}_{t}$ proceeds analogously. Thus the family $\Sigma_{t}$ is totally geodesic within $V$ since its second fundamental form is given by ${\cal K}=\chi^{+}_{t}-\chi^{-}_{t}$.
@@
1.) $\mathbf{H_{t}\neq 0\leavevmode\nobreak\ for\leavevmode\nobreak\ some\leavevmode\nobreak\ t\neq 0}$.
@@
$\Sigma_{y}$$\Sigma_{x}$$\Sigma$$\Sigma_{s}$$\nu$$\Sigma_{s,u}$$l^{-}$xxyy
@@
Therefore, from condition (1.2) of Theorem 2, $\operatorname{tr}_{\Sigma_{s}}{\cal K}\geq 0$, and thus (1.1) implies that $\theta^{-}_{s}<0$. As in the proof of Lemma 1 this property is inherited by the cover, i.e. $\tilde{\theta}^{-}_{s}<0$, and application of Penrose’s singularity theorem (with respect to the non-compact end in the $\nu -$direction from $\tilde{\Sigma}$) contradicts the required null geodesic completeness within $M$.
2.) $\mathbf{H_{t}\equiv 0\leavevmode\nobreak\ \forall\leavevmode\nobreak\ t}$.
@@
To show the latter we recall the relations $\chi=2(1-\mathfrak{g})$ and $\chi=2-\mathfrak{g}$ between genus $\mathfrak{g}$ and Euler number $\chi$ for orientable and non-orientable 2-surfaces, respectively. The Euler number is also the sum of all indices of the (isolated) zeros of $\xi$, by virtue of the Poincaré-Hopf theorem. The zeros of $\xi$ are necessarily isolated since $\Sigma$ is two-dimensional [10, Thm. 8.1.5], and the index of an isolated zero of a Killing vector field is +1 since they’re rotations in a neighborhood of the zero. This restricts the topology of an orientable surface $\Sigma$ (or of its orientable double cover $\tilde{\Sigma}$) to the sphere or the torus, and to the projective plane or the Klein bottle for a non-orientable $\Sigma$.
@@ 3 Proofs of the main results
We claim that $V$ or a finite cover thereof has nontrivial second homology. To see this, recognize that by the prime decomposition of three-manifolds, $V$ is homeomorphic to $V_{1}\#\dotsb\#V_{k}$, where each $V_{i}$ is a prime manifold. Consequently, $H_{2}(V)=H_{2}(V_{1})\oplus\dotsb\oplus H_{2}(V_{k})$. If some $V_{i}$ is topologically $\mathbb{S}^{1}\times\mathbb{S}^{2}$, then $H_{2}(V)\neq 0$. If some $V_{i}$ is aspherical (i.e., the universal cover of $V_{i}$ is contractible), then $V_{i}$ has a finite cover with nonvanishing second homology by the positive resolution of the virtual $b_{1}>0$ conjecture [5]; in this case, $V$ also has a finite cover with nontrivial second homology (see [2, Lem. 6]). Lastly, if each $V_{i}$ is a spherical space, then the arguments below the proof of [2, Lem. 6] show the desired result. This proves the claim.
@@
We first prove Theorem 1 by showing that conclusion (iii) holds. Since $\Sigma$ is two-sided and nonseparating, we can cut $V$ along $\Sigma$ to produce a compact manifold $W$ with two boundary components – each one isometric to $\Sigma$. We construct a covering $\tilde{V}$ of $V$ by gluing $\mathbb{Z}$ copies of this compact manifold $W$ end to end. See e.g. the proof of Prop. 5 of [2]. It’s clear that the hypotheses of Lemma 1 are satisfied for some particular copy of $\Sigma$ in the cover $\tilde{V}$. Thus there is a neighborhood of $\Sigma$ within $V$, diffeomorphic to $(-\epsilon,\epsilon)\times\Sigma$, such that its leaves $\Sigma_{t}$ are totally geodesic for all $t$. Using the traced Gauss equation on each $\Sigma_{t}$, the compactness of $V$ implies that the scalar curvature of the $\Sigma_{t}$ stays uniformly bounded. Therefore, by standard results (e.g. the "Compactness Theorem" in Sect. 2 of [13]), the foliation has a limit surface $\Sigma_{\epsilon}$, which is embedded, compact, and minimal. Moreover $\Sigma_{\epsilon}$ is two-sided: if not, then $\Sigma$ would have separated $V$. Therefore, by applying Lemma 1 again, there is a totally geodesic foliation near $\Sigma_{\epsilon}$. The maximum principle implies that $\Sigma_{\epsilon}$ is diffeomorphic to $\Sigma$. A continuity argument now shows that $W\approx[0,1]\times\Sigma$ and each $\Sigma_{t}$ is totally geodesic. Thus $V$ is a surface bundle over $\mathbb{S}^{1}$ with totally geodesic fibers.
@@
By general covering space theory, there is a covering $p\colon\tilde{V}\to V$ such that $\pi_{1}(\tilde{V})\cong\pi_{1}(\Sigma)$. By the lifting criterion, $f$ lifts to $F\colon\Sigma\to\tilde{V}$ which is a two-sided minimal embedding (in fact it is least area in its homotopy class by Theorem 3.4 in [15]). It must be that $\Sigma$ separates $\tilde{V}$. If not, then there is a loop in $\tilde{V}$ which traverses $\Sigma$ once and therefore generates an element in $\pi_{1}(\tilde{V})\setminus\pi_{1}(\Sigma)$, contradicting the construction of $\tilde{V}$ [^2] This is a consequence of a more general fact. In short: If $\Sigma$ were nonseparating in $\tilde{V}$, then $\pi_{1}(\tilde{V})\cong A*_{\pi_{1}(\Sigma)}$ (see e.g. [16, Proposition 1.2]) but then $\lvert\pi_{1}(\tilde{V}):\pi_{1}(\Sigma)\rvert$$\infty$, a contradiction to our assumptions..
@@
Thus, there are nonorientable limit surfaces $\Sigma_{1}$ and $\Sigma_{2}$. By gluing two copies of $V\setminus(\Sigma_{1}\cup\Sigma_{2})$ along their two boundary components, we get a double cover $V^{\prime}$ of $V$. The boundary components of $V\setminus(\Sigma_{1}\cup\Sigma_{2})$ are diffeomorphic to $\Sigma$ (this can be seen via [20, Thm. 10.5] or another maximum principle argument within $V^{\prime}$). Thus, $V^{\prime}$ is a surface bundle over $\mathbb{S}^{1}$ with fiber $\Sigma$. Hence $\Sigma_{1}$, $\Sigma_{2}$ are double covered by $\Sigma$. Subsequently, $\Sigma_{1}\cong\Sigma_{2}$ by the classification theorem of closed surfaces. Therefore, by construction, $V$ has a semibundle structure, meaning it is the union of two twisted $I$-bundles over $\Sigma_{1}\cong\Sigma_{2}$ glued along their common boundary $\Sigma$. That is, $V=V_{1}\cup V_{2}$ where $V_{i}\cong\Sigma_{i}\tilde{\times}I$.

### v3

@@ Abstract
Improving a singularity theorem in General Relativity by Galloway and Ling we show the following (cf. Theorem 1): If a globally hyperbolic spacetime $M$ satisfying the null energy condition contains a closed, spacelike Cauchy surface $(V,g,\cal K)$ (with metric $g$ and extrinsic curvature $\cal K$}) which is 2-convex (meaning that the sum of the lowest two eigenvalues of $\cal K$} is non-negative), then either $M$ is past null geodesically incomplete, or $V$ is a spherical space, or $V$ or some finite cover $\tilde{V}$ is a surface bundle over the circle, with totally geodesic fibers. Moreover, (cf. Theorem 2) if $(V,g,{\cal K})$ admits a $U(1)$ isometry group with corresponding Killing vector $\xi$, we can relax the convexity requirement in terms of a decomposition of $\cal K$} with respect to the directions parallel and orthogonal to $\xi$. Finally, (cf. Propositions 1-3) in the special cases that $V$ is either nonorientable, or non-prime, or an orientable Haken manifold with vanishing second homology, we obtain stronger statements in both Theorems without passing to covers. While our results do not use Einstein’s equations, we provide several classes of $\Lambda$-vacuum solutions of these equations as examples.
@@ 1 Introduction
The classic singularity theorems due to Hawking and Penrose [24] yield geodesic incompleteness in different settings, namely "cosmological" and "black hole" ones. In the former setting one requires a compact Cauchy surface and a condition on its second fundamental form, while the latter setting assumes noncompact data containing a trapped or marginally trapped surface. In either case, convergence ("energy-") conditions on the spacetime Ricci tensor $\mathrm{Ric}$ are also necessary. In the sequel we will only need the null energy condition (NEC), namely $\mathrm{Ric}(X,X)\geq 0$ for all null vectors $X$.
@@
    Assume $V$ is $2$-convex, i.e. the sum of the smallest two eigenvalues of the future second fundamental form $\mathcal{K}$ is non-negative.
@@
We conclude here with sketching the strategy of our proofs. For simplification, we assume that $V$ is not a spherical space, and we pass to a finite, orientable cover $\tilde{V}$ whenever necessary or useful. Our key prerequisite is the prime decomposition of orientable 3-manifolds, see (3.1) below. Moreover, from the positive resolution of the virtual positive $b_{1}$-conjecture [1], it follows that $V$ or a finite cover $\tilde{V}$ has non-vanishing second homology $H_{2}$. As is well-known, this implies that $V$ (or $\tilde{V}$) contains an embedded, non-separating, two-sided minimal surface $\Sigma$ of least area. By a known Lemma (Lemma 0), it follows that a neighborhood of $\Sigma$ can be foliated by CMC surfaces $\Sigma_{t}$. To prove Theorem 1 we now invoke the convexity requirement on $\cal K$} which, via (1.1), yields control on the signs of the null expansions $\theta^{\pm}$ on each $\Sigma_{t}$. Together with the null energy condition, this allows an application of Penrose’s singularity theorem on a suitable infinite cover of $V$ (or $\tilde{V}$). This gives past null geodesic incompleteness except for the degenerate cases stated in (ii), (iii) of Theorem 1. In case (iii) compactness theorems allow us to extend the local folitaion of Lemma 1 to all of $V$.
@@
We remark that Theorem 1 can be proven in a different way, namely by using the positive resolution of the virtual Haken conjecture and Proposition 1. In fact, this detour was taken by the authors originally. We are grateful to Greg Galloway for bringing to our attention the positive resolution of the virtual positive $b_{1}$ conjecture which greatly simplifies the proof. A similar simplification also applies to the proof of the main result in [16], reformulated as Theorem 0 above (see [17]).
@@ 2 Preliminaries
The following lemma is a standard application of the inverse function theorem. For a proof of the more involved $\lambda=0$ case, see the proof of Thm. 2.38 in [29] or [18, Appendix A].
@@
Let $\Sigma$ be an embedded stable minimal surface within a Riemannian manifold $V$. Then there is a neighborhood $U$ of $\Sigma$ within $V$, diffeomorphic to $(-\epsilon,\epsilon)\times\penalty 10000\ \Sigma$, such that each leaf $\Sigma_{t}:=\{t\}\times\Sigma$ has constant mean curvature.
@@
Proof. The following facts will be used to prove this Lemma as well as Lemma 2. First, there is a spacetime cover $P\colon\tilde{M}\to M$ such that $p=P|_{\tilde{V}}$, $\tilde{V}$ is a Cauchy surface for $\tilde{M}$, and the second fundamental form on $\tilde{V}$ is the pullback $\tilde{\cal K}=P^{*}\cal K$, see [16, Lem. 4]. Moreover, the null convergence condition also “lifts”, and $M$ is null geodesically complete if and only if $\tilde{M}$ is. Lastly, if a surface $\Sigma^{\prime}$ embedded in $V$ is constructed via a foliation from $\Sigma$ within $V$, then $\Sigma^{\prime}$ lifts to some $\tilde{\Sigma}^{\prime}$, which, like $\tilde{\Sigma}$, also separates $\tilde{V}$ into two noncompact ends.
@@
2.) $\mathbf{H_{t}\equiv 0\penalty 10000\ \forall\penalty 10000\ t}$.
@@
say $\tilde{W}_{t}^{+}$, does not vanish identically on $\tilde{\Sigma}_{t}$. By virtue of Raychaudhuri’s equation, $\frac{d}{ds}\tilde{\theta}_{t}^{+}=-\tilde{W}_{t}^{+},$ and another application of Lemma 5.2 in [6] we find surfaces $\tilde{\Sigma}^{\prime}$ near the past light cone of $\tilde{\Sigma}_{t}$ with $\tilde{\theta}^{\prime+}<0$. Another application of Penrose’s theorem in the cover $\tilde{M}$ and projecting down to $M$ contradicts past null geodesic completeness within $M$. Therefore both $(+)$-terms on the r.h.s. of (2.2) have to vanish individually. Clearly, the case $\tilde{W}^{-}_{t}$ proceeds analogously. Thus the family $\Sigma_{t}$ is totally geodesic within $V$ since its second fundamental form is given by $\chi^{+}_{t}-\chi^{-}_{t}$.
@@
1.) $\mathbf{H_{t}\neq 0\penalty 10000\ for\penalty 10000\ some\penalty 10000\ t\neq 0}$.
@@
$\Sigma_{y}$$\Sigma_{x}$$\Sigma$$\Sigma_{s}$$\nu$$\Sigma_{s,u}\ell^{-}$xxyy
@@
Therefore, from condition (1.2) of Theorem 2, $\operatorname{tr}_{\Sigma_{s}}{\cal K}\geq 0$, and thus (1.1) implies that $\theta^{-}_{s}<0$. As in the proof of Lemma 1 this property is inherited by the cover, i.e. $\tilde{\theta}^{-}_{s}<0$, and application of Penrose’s singularity theorem (with respect to the noncompact end in the $\nu -$direction from $\tilde{\Sigma}$) contradicts the required null geodesic completeness within $M$.
2.) $\mathbf{H_{t}\equiv 0\penalty 10000\ \forall\penalty 10000\ t}$.
@@
To show the latter we recall the relations $\chi=2(1-\mathfrak{g})$ and $\chi=2-\mathfrak{g}$ between genus $\mathfrak{g}$ and Euler number $\chi$ for orientable and nonorientable 2-surfaces, respectively. The Euler number is also the sum of all indices of the (isolated) zeros of $\xi$, by virtue of the Poincaré-Hopf theorem. The zeros of $\xi$ are necessarily isolated since $\Sigma$ is two-dimensional [33, Thm. 8.1.5], and the index of an isolated zero of a Killing vector field is +1 since they’re rotations in a neighborhood of the zero. This restricts the topology of an orientable surface $\Sigma$ (or of its orientable double cover $\tilde{\Sigma}$) to the sphere or the torus, and to the projective plane or the Klein bottle for a nonorientable $\Sigma$.
@@ 3 Proofs of the main results
We claim that $V$ or a finite cover thereof has non-trivial second homology. To see this, recognize that by the prime decomposition of three-manifolds, $V$ is homeomorphic to $V_{1}\#\dotsb\#V_{k}$, where each $V_{i}$ is a prime manifold. Consequently, $H_{2}(V)=H_{2}(V_{1})\oplus\dotsb\oplus H_{2}(V_{k})$. If some $V_{i}$ is topologically $\mathbb{S}^{1}\times\mathbb{S}^{2}$, then $H_{2}(V)\neq 0$. If some $V_{i}$ is aspherical (i.e., the universal cover of $V_{i}$ is contractible), then $V_{i}$ has a finite cover with nonvanishing second homology by the positive resolution of the virtual $b_{1}>0$ conjecture [1]; in this case, $V$ also has a finite cover with non-trivial second homology (see [16, Lem. 6]). Lastly, if each $V_{i}$ is a spherical space, then the arguments below the proof of [16, Lem. 6] show the desired result. This proves the claim.
@@
We first prove Theorem 1 by showing that conclusion (iii) holds. Since $\Sigma$ is two-sided and nonseparating, we can cut $V$ along $\Sigma$ to produce a compact manifold $W$ with two boundary components – each one isometric to $\Sigma$. We construct a covering $\tilde{V}$ of $V$ by gluing $\mathbb{Z}$ copies of this compact manifold $W$ end to end. See e.g. the proof of Prop. 5 of [16]. It’s clear that the hypotheses of Lemma 1 are satisfied for some particular copy of $\Sigma$ in the cover $\tilde{V}$. Thus there is a neighborhood of $\Sigma$ within $V$, diffeomorphic to $(-\epsilon,\epsilon)\times\Sigma$, such that its leaves $\Sigma_{t}$ are totally geodesic for all $t$. Using the traced Gauss equation [29, Cor. 2.7] on each $\Sigma_{t}$, the compactness of $V$ implies that the scalar curvature of the $\Sigma_{t}$ stays uniformly bounded. Therefore, by standard results (e.g. the "Compactness Theorem" in Sect. 2 of [3]), the foliation has a limit surface $\Sigma_{\epsilon}$, which is embedded, compact, and minimal. Moreover $\Sigma_{\epsilon}$ is two-sided: if not, then $\Sigma$ would have separated $V$. Therefore, by applying Lemma 1 again, there is a totally geodesic foliation near $\Sigma_{\epsilon}$. The maximum principle implies that $\Sigma_{\epsilon}$ is diffeomorphic to $\Sigma$. A continuity argument now shows that $W\approx[0,1]\times\Sigma$ and each $\Sigma_{t}$ is totally geodesic. Thus $V$ is a surface bundle over $\mathbb{S}^{1}$ with totally geodesic fibers.
@@
By general covering space theory, there is a covering $p\colon\tilde{V}\to V$ such that $\pi_{1}(\tilde{V})\cong\pi_{1}(\Sigma)$. By the lifting criterion, $f$ lifts to $F\colon\Sigma\to\tilde{V}$ which is a two-sided minimal embedding (in fact it is least area in its homotopy class by Theorem 3.4 in [21]). It must be that $\Sigma$ separates $\tilde{V}$. If not, then there is a loop in $\tilde{V}$ which traverses $\Sigma$ once and therefore generates an element in $\pi_{1}(\tilde{V})\setminus\pi_{1}(\Sigma)$, contradicting the construction of $\tilde{V}$ [^2] This is a consequence of a more general fact. In short: If $\Sigma$ were nonseparating in $\tilde{V}$, then $\pi_{1}(\tilde{V})\cong A*_{\pi_{1}(\Sigma)}$ (see e.g. [40, Proposition 1.2]) but then $\lvert\pi_{1}(\tilde{V}):\pi_{1}(\Sigma)\rvert$$\infty$, a contradiction to our assumptions..
@@
Thus, there are nonorientable limit surfaces $\Sigma_{1}$ and $\Sigma_{2}$. By gluing two copies of $V\setminus(\Sigma_{1}\cup\Sigma_{2})$ along their two boundary components, we get a double cover $V^{\prime}$ of $V$. The boundary components of $V\setminus(\Sigma_{1}\cup\Sigma_{2})$ are diffeomorphic to $\Sigma$ (this can be seen via [26, Thm. 10.5] or another maximum principle argument within $V^{\prime}$). Thus, $V^{\prime}$ is a surface bundle over $\mathbb{S}^{1}$ with fiber $\Sigma$. Hence $\Sigma_{1}$, $\Sigma_{2}$ are double covered by $\Sigma$. Subsequently, $\Sigma_{1}\cong\Sigma_{2}$ by the classification theorem of closed surfaces. Therefore, by construction, $V$ has a semibundle structure, meaning it is the union of two twisted $I$-bundles over $\Sigma_{1}\cong\Sigma_{2}$ glued along their common boundary $\Sigma$. That is, $V=V_{1}\cup V_{2}$ where $V_{i}\cong\Sigma_{i}\tilde{\times}I$.